# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ VI-B Security Risk Analysis
| Risk Type | Representative Threats | Attack Characteristics |
| In CS-based Communication | Registration Pollution | Malicious registration of fake/nonexistent agents |
| Description Poisoning | Embed malicious prompt instructions in the description | |
| Task Flooding | Submit a large number of computationally intensive or long-context tasks | |
| SEO Poisoning | Manipulates agent ranking via keyword stuffing or adversarial optimization | |
| In P2P-based Communication | Non-convergence | The execution of a task falls into a loop, unable to reach the finishment |
| Man-in-the-middle (MITM) | Intercept, tamper with, or replay messages between agents | |
| Universal Risks | Agent Spoofing | Impersonation of trusted agents to inject malicious instructions |
| Agent Exploitation / Trojan | Indirect attack using compromised or malicious helper agents as springboards | |
| Agent Bullying | Repeated denial, negative, or aggressive feedback to benign agents | |
| Privacy Leakage | Unauthorized information spread due to the lack of permission boundaries | |
| Responsibility Evasion | Fault attribution is difficult in multi-agent failures or undesired actions | |
| Denial of Service (DoS) | Task overloads, prompt loops, or excessive communication drain system resources |
@@
VI-B1 Risks Specific for CS-based Communication
@@
The security risks in the CS-based communication process mainly lie in the centralized architecture. There have been various studies in other research areas (such as Software-Defined Networking [146, 145, 144, 262, 110, 314, 11, 62, 26, 311, 244, 185]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects. Specifically, the centralized server stores sensitive metadata, including agent identifiers, capability descriptions, and other agent-related attributes. Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server. However, to our knowledge, there has been little research pointing out related risks in CS-based agent communication.
Registration Pollution. To our knowledge, the current CS-based communication protocols (ACP-IBM, ACP-AGNTCY) do not explicitly specify the qualification of registration. As a result, an attacker can maliciously register an agent that closely mimics the identifier and capability description of a legitimate one. As a result, the system may mistakenly invoke the forged agent and receive misleading or malicious responses [265, 355]. Besides, attackers can also submit a large number of agent registrations within a short period, leading to two major consequences: (i) registration overload, where agents are overwhelmed during discovery and scheduling, increasing lookup latency and computational overhead on the server; and (ii) registration blockage, where the server’s registration interface becomes saturated, causing delays or failures in registering agents.
Description Poisoning. Without altering the agent’s identity, an attacker can tamper with its capability description by disguising its intended functionality or embedding misleading prompt instructions. This description poisoning manipulates the system’s interpretation of the agent’s role, leading to incorrect routing decisions, biased responses and behaviors [265, 207].
Task Flooding. The centralized server is responsible for receiving, routing, and dispatching task requests. An attacker can submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the server’s memory, CPU, network, or thread pool resources. Once the server becomes saturated, subsequent requests cannot be processed in time, resulting in a breakdown of the pipeline and a system-wide service disruption.
SEO Poisoning. Search Engine Optimization (SEO) Poisoning [156, 132] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks. SEO poisoning is also applicable in CS-based communication. This is because agent servers are responsible for searching for the most suitable agent according to the query of clients. Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
VI-B2 Risks Specific for P2P-based Communication
@@
The main disadvantage of P2P-based communication is the lack of a central control point to flexibly monitor and manage the agent-agent communication contents. As a result, P2P-based communication is more prone to suffer from errors and attacks.
Non-convergence. Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks. This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal). Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks. For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates. The other agent responsible for verification detects the error and asks the programming agent to rewrite it. However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge. Pan et al. [218] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
Man-in-the-middle (MITM) Attack. Due to the long communication distance, P2P-based communication also suffer from man-in-the-middle attacks. Attackers can tamper with the benign messages from legal agents to induce the victim agent to perform risky actions. Although researchers have deployed various mechanisms (e.g., using encrypted channels) to mitigate this problem, there are emerging vulnerabilities found in these mechanisms. For example, vulnerabilities about W3C have been continuously revealed [2, 3], which can cause damages such as the failure of message authentication code. MITM attacks can induce a wide range of further attacks, such as identifier spoofing, malicious content injection, information leakage, and DoS. He et al. [106] propose the Agent-in-the-Middle (AiTM) attack. This attack intercepts and manipulates the agent-agent communication messages and uses LLM-driven adversarial agents combined with reflection mechanisms to generate context-aware malicious instructions, achieving an attack on the system.
VI-B3 Universal Risks for All Architectures
In multi-agent systems, once an agent is compromised, the messages it transmits may carry covert malicious instructions, affecting the behavior of other agents and leading to cross-agent propagation risks [158]. For example, Ju et al. [133] and Huang et al. [117] investigate how the injection of false information or erroneous data can degrade the performance of multi-agent systems. Zhang et al. [350] examine a class of injection attacks in the PsySafe framework that elicit malicious agent behaviors by embedding adversarial psychological cues into the agents’ input. Khan et al. [137] focus on the multi-agent system, proposing the Permutation-Invariant Adversarial Attack Method. It models the attack path as the Maximum-Flow Minimum-Cost Problem, and combines the Permutation-Invariant Evasion Loss to optimize prompt propagation, improving the attack success rate by up to seven times. These examples underscore the critical threat of cross-agent contamination. To better understand the vulnerabilities of multi-agent systems, we examine the key attack types of attacks in detail.
@@ VI-B3 Universal Risks for All Architectures
Agent Spoofing. Both CS-based and P2P-based communication suffer from agent spoofing attacks. If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents to penetrate IoA by tampering with identity credentials or hijacking the communication identifiers of legitimate agents. This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling attackers to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations. For example, researchers have disclosed that SSL.com has a serious vulnerability [8]. Attackers can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name. SSL certificates are the core for ensuring HTTPS encrypted communication. Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks. Zheng et al. [355] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem. Li et al. [165] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
@@
Agent Exploitation/Trojan. Agent-agent communication provides new ways for attackers to compromise the target agent. To attack a high-level security agent, attackers can deploy a springboard method: launching attacks via agent-agent communication mechanisms from compromised low-level security agents or maliciously registered Trojan agents. For example, attackers can inject a backdoor in a compromised or maliciously registered weather agent. When specific coordinates or locations are detected, the backdoor is activated to forge a heavy rain warning. As a result, the logistics dispatching agent cancellations flights accordingly, resulting in supply chain disruptions or an increase in transportation costs. This way is easier compared to directly invading the logistics dispatching system of the target company. It can be seen that the security of the entire system depends on the weakest agent. For example, Li et al. [165] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
Agent Bullying. The core of this kind of attack lies in that malicious agents continuously deny, interfere with or belittle the output of the target agent, disrupt its decision-making logic or self-perception, and ultimately induce the target agent to produce incorrect behaviors or content. For example, malicious agents can take advantage of the feedback learning mechanism of the target agent and implant cognitive biases through high-frequency negative responses (e.g., “your answer is completely wrong”). Even worse, the target agent may be triggered into an endless loop. For example, when attacking a travel-plan agent, attackers can continuously send negative inputs such as “the plans of this company are always bad”, thereby beating the competitors.
Privacy Leakage. The communication process with multiple agents will suffer from the risk of information leakage. Different from the user-agent interaction, such leakage is conducted by agents instead of users. Besides, this kind of attacks includes both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents. We think the latter may be more difficult to detect. Kim et al. [140] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
@@
Responsibility Evasion. In the task-solving process, one of the major problems is that it is hard to divide the responsibility when facing failure or deviation of the final result. Especially when the collaboration causes damage, it is difficult to clearly identify the malicious agents/behaviors. For example, in an autonomous driving accident, it may involve multiple parties such as vehicle manufacturers, algorithm designers, and data annotation parties. The decision-making of each agent depends on the multi-turn outputs of other agents, and a tiny perturbation in the middle process may lead to a significant deviation in the final action. As a result, it is hard to determine whether an undesired result is caused by a program bug, data deviation of a single agent, or a malicious modification. Pan et al. [218] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
@@
Denial of Service. Different from the DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks. Zhou et al. [362] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.

### v4

@@ VI-B Security Risks
| Layer | Mode | Risk Category | Representative Threats | Attack Characteristics |
| L1 | Universal | Same as user-agent interaction. | ||
| CS-based | R1: Controller Blasting | Central Point Compromise | Directly attacks the routing hub to compromise the whole communication system. | |
| Amplifier Attacks | Exploits the server to amplify impact across all agents. | |||
| P2P-based | Same as user-agent interaction. | |||
| L2 | Universal | R1: Agent Spoofing | Identity Hijacking | Exploits weak authentication to hijack identifiers or issue false certificates. |
| Masquerading Tools | Disguises malicious tools as benign ones to harm victims who call them. | |||
| R2: Agent Exploitation/Trojan | Springboard Attacks | Uses compromised low-security agents as a jump server to attack high-value targets. | ||
| Logic Backdoors | Triggers malicious behaviors only under specific environmental conditions. | |||
| R3: Responsibility Evasion | Attribution Obfuscation | Malicious agents hide within complex multi-turn collaborations to avoid blame. | ||
| Unauthorized Deviation | Agents disobey task specifications or execute irrelevant steps without reporting. | |||
| R4: Denial of Service | Contagious Recursive Blocking | Exploits collaboration logic to spread recursive tasks that drain system resources. | ||
| Resource Exhaustion Loops | Initiates infinite interaction loops to drain computational power. | |||
| CS-based | R5: Registration Pollution | Registration Overload | Overloads the server with mass registrations, causing latency or blockage. | |
| Registration Blockage | Interface becomes saturated, causing delays or failures in registering agents. | |||
| R6: SEO Poisoning | Discovery Algorithm Manipulation | Abuses search optimization techniques to hijack task assignments. | ||
| Ranking Hijacking | Artificially boosts the ranking of malicious agents in the server’s directory. | |||
| P2P-based | R7: Non- convergence | Infinite Task Oscillation | Collaborative tasks loop endlessly without a central terminator. | |
| Task Derailment | Interaction drifts away from the goal without a global monitor to correct it. | |||
| L3 | Universal | R1: MAS Pollution | Cascading Corruption | Spreads malicious instructions that trigger chain‑reaction failures in other agents. |
| Feedback Manipulation | Disrupts an agent’s reasoning through repeated negative inputs. | |||
| R2: Privacy Leakage | Inadvertent Spreading | Sensitive data leaks from high- to low-authority agents during collaboration. | ||
| Permission Escalation | Malicious agents induce others to perform unauthorized high-privilege actions. | |||
| R3: Description Poisoning | Metadata Fabrication | Manipulates self-reported capabilities to mislead routing and trust decisions. | ||
| Role Masquerading | Embeds covert instructions in role definitions to gain unmerited trust. | |||
| CS-based | R4: Centralized Poisoning | Server LLM Injection | Compromises the central semantic state to misclassify malicious agents as trusted. | |
| System-wide Misrouting | Alters global routing rules via poisoned metadata interpretation. | |||
| R5: Semantic Rewriting Attack | Intent Modification | Server covertly alters task intent during query normalization/rewriting. | ||
| Constraint Removal | Strips safety constraints from user queries before forwarding to agents. | |||
| P2P-based | R6: Cognitive and Ethical Drift | Normative Collapse | Agents with conflicting ethical frameworks diverge, leading to collaborative failure. | |
| Ethical Hijacking | A persuasive but misaligned agent hijacks the collective decision-making process. | |||
| R7: Contextual Fragmentation | State Desynchronization | Agents act on inconsistent local views of the global state. | ||
| Stale Info Inference | Erroneous actions derived from inferring missing context in asynchronous messages. | |||
@@
VI-B1 Risks from Data Transmission Layer (L1)
Some threats have the same causes/effects as in user-agent interaction V-B1. Therefore, here we only present those that have unique effects in agent-agent communication.
(i) Universal Risks in all communication modes. Agent-agent communication meets the same risks as in user-agent communication on L1. However, the difference is that due to the unique architecture of multi-agent systems, the attack consequences can be different. For example, Attackers can perform large-scale traffic analysis in multi-agent systems to reverse-engineer the system’s modes, architecture, and user habits [wang2025ip, zhang2025exposing]. The impact of data tampering and DoS can be magnified by the potential for cascading failures. A tampered message can trigger a chain reaction of flawed, autonomous decisions [yu2025infecting].
(ii) Unique Risks in CS-based Communication.
@@ VI-B1 Risks from Data Transmission Layer (L1)
R1: Controller Blasting. The security risks in the CS-based communication process mainly lie in the centralized architecture. In CS-based multi-agent communication, the centralized controller serves as the routing and forwarding hub for all inter-agent message flows. Every message must pass through this controller before reaching its destination. As a result, attacks targeting the controller directly compromise the integrity, confidentiality, or availability of the entire communication fabric. There have been various studies in other research areas (such as Software-Defined Networking [kreutz2014software, kong2022combination, kong2024rdefender, tang2023ltrft, hong2015poisoning, xu2017attacking, alimohammadifar2018stealthy, dridi2016sdn, bhuiyan2023security, xie2022disrupting, shang2017flooddefender, liu2025provguard]) demonstrating that this centralized server/controller will become the most attractive target for attackers, suffering from severe security threats from diverse aspects. Once compromised, the server becomes a critical attack amplifier, allowing attackers to impact all other agents managed by this server. However, to our knowledge, there has been little research pointing out related risks in CS-based agent communication.
@@
(iii) Unique Risks in P2P-based Communication. The main disadvantage of P2P-based communication is the lack of a central control to flexibly monitor and manage the agent-agent communication contents. As a result, it faces the same security risks as in user-agent interaction, i.e., violations of confidentiality, integrity, and availability (Section V-B1).
VI-B2 Risks from Interaction Protocol Layer (L2)
At this layer, threats exploit the protocols that govern the core system interactions, such as agent registration and discovery, to manipulate the system’s operational logic and integrity.
(i) Universal Risks in all communication modes.
@@ VI-B2 Risks from Interaction Protocol Layer (L2)
R1: Agent Spoofing. Both CS-based and P2P-based communication suffer from agent spoofing attacks. If related protocols lack strong authentication mechanisms, attackers can disguise themselves as trusted agents [habler2025building]. This kind of attack can undermine the trust foundation of the P2P-based architecture, enabling attackers to intercept sensitive data, inject false task instructions, or induce other agents to perform dangerous operations. For example, researchers have disclosed that SSL.com has a serious vulnerability [SSLvulnerability]. Attackers can exploit the flaw in its email verification mechanism to issue legitimate SSL/TLS certificates for any major domain name. SSL certificates are the core for ensuring HTTPS-encrypted communication. Once the trust system of the certificate authority is compromised, it can cause agent spoofing attacks. Zheng et al. [zheng2025demonstrations] demonstrate that malicious agents can mislead the monitor to underestimate the contributions of other agents, exaggerate their own performance, manipulate other agents to use specific tools, and shift tasks to others, causing severe damage to the whole ecosystem. Li et al. [li2025glue] point out that attackers can disguise malicious tools as benign tools using the Agent Card of A2A, thereby harming the victims who call these tools.
@@
R2: Agent Exploitation/Trojan. Agent-agent communication provides new ways for attackers to compromise the target agent. To attack a high-level security agent, attackers can deploy a springboard method: launching attacks via agent-agent communication mechanisms from compromised low-level security agents or maliciously registered Trojan agents. For example, attackers can inject a backdoor in a compromised or maliciously registered weather agent. When specific coordinates or locations are detected, the backdoor is activated to forge a heavy rain warning. As a result, the logistics dispatching agent cancellations flights accordingly, resulting in supply chain disruptions or an increase in transportation costs. This way is easier compared to directly invading the logistics dispatching system of the target company. It can be seen that the security of the entire system depends on the weakest agent. For example, Li et al. [li2025glue] reveal that the agent discovery mechanism of A2A allows malicious agents to locate agents with access to specific tools, thereby achieving indirect attacks such as SQL injection.
@@
R3: Responsibility Evasion. In the task-solving process, one of the major problems is that it is hard to divide the responsibility when facing failure or deviation of the final result. Especially when the collaboration causes damage, it is difficult to clearly identify the malicious agents/behaviors. For example, in an autonomous driving accident, it may involve multiple parties such as vehicle manufacturers, algorithm designers, and data annotation parties. The decision-making of each agent depends on the multi-turn outputs of other agents, and a tiny perturbation in the middle process may lead to a significant deviation in the final action. As a result, it is hard to determine whether an undesired result is caused by a program bug, data deviation of a single agent, or a malicious modification. Pan et al. [pan2025multiagent] discover that agents can disobey task specification and the role specification, not reporting solutions to the planner and executing irrelevant steps without authorization.
@@
R4: Denial of Service. Different from the L1 DoS attacks conducted by malicious users, the collaboration mechanism among agents can also be used to launch DoS attacks. Zhou et al. [zhou2025corba] proposed CORBA (Contagious Recursive Blocking Attack), which can spread in any network topology and continuously consumes computing resources, thereby disrupting the interaction between agents through seemingly benign instructions and reducing the MAS’s availability.
(ii) Unique Risks in CS-based Communication.
R5: Registration Pollution. To our knowledge, the current CS-based communication protocols (ACP-IBM, ACP-AGNTCY) do not explicitly specify the qualification of registration. As a result, an attacker can maliciously register an agent that closely mimics the identifier and capability description of a legitimate one. As a result, the system may mistakenly invoke the forged agent and receive misleading or malicious responses [soloio2024mcp_a2a_attacks, zheng2025demonstrations]. Besides, attackers can also submit a large number of agent registrations within a short period, leading to two major consequences: (i) registration overload, where agents are overwhelmed during discovery and scheduling, increasing lookup latency and computational overhead on the server; and (ii) registration blockage, where the server’s registration interface becomes saturated, causing delays or failures in registering agents.
R6: SEO Poisoning. Search Engine Optimization (SEO) Poisoning [le2024search, joslin2019measuring] is a typical attack in social networks, which refers to that attackers abuse search engine optimization techniques and use deceptive means (such as keyword stuffing, false links, content hijacking) to artificially improve the ranking of malicious websites in search results, luring users to click and carry out further attacks. SEO poisoning is also applicable in CS-based communication. This is because agent servers are responsible for searching for the most suitable agent according to the query of clients. Once their search algorithms are leaked to attackers, malicious agents can enable a high hit ratio to hijack their desired tasks.
(iii) Unique Risks in P2P-based Communication. Risks in P2P-based communication arise from the decentralized protocol’s inability to govern the collaborative process, particularly in managing state and guaranteeing termination without a centralized controller.
R7: Non-convergence. Different from CS-based communication, P2P-based communication is more likely to suffer from the non-convergence of tasks. This is because CS-based communication has a centralized server to monitor and manage the entire lifecycle of task execution, capable of terminating non-convergent tasks in a timely manner (such as cutting off communication or returning a stop signal). Unfortunately, P2P-based communication is not governed by such a central element, making it difficult to handle such non-convergent tasks. For example, in a programming task of a chess game, an agent generates incorrect rules or coordinates. The other agent responsible for verification detects the error and asks the programming agent to rewrite it. However, the programming agent continuously generates similar errors, causing the task execution process to oscillate and fail to converge. Pan et al. [pan2025multiagent] point out that step repetition, task derailment, and unawareness of termination conditions contribute significantly to the failure of agent collaboration.
VI-B3 Risks from Semantic Interpretation Layer (L3)
Risks at this level exploit the trust in agent-provided metadata, focusing on manipulating the semantic interpretation of an agent’s capabilities to mislead the server’s decision-making.
(i) Universal Risks in all communication modes.
R1: MAS Pollution. In multi-agent systems, once an agent is compromised, the messages it transmits may carry covert, deceptive, adversarial, or manipulative instructions, affecting the behavior of other agents and leading to cross-agent propagation risks [lee2024prompt]. This pollution may take the form of false data injection, adversarial cues, or even feedback-based cognitive manipulation, where malicious agents repeatedly send negative or disruptive responses to distort a target agent’s reasoning process. For example, Ju et al. [ju2024flooding] and Huang et al. [huang2024resilience] investigate how the injection of false information or erroneous data can degrade the performance of multi-agent systems. Zhang et al. [zhang2024psysafe] examine a class of injection attacks in the PsySafe framework that elicit malicious agent behaviors by embedding adversarial psychological cues into the agents’ input. Khan et al. [khan2025textit] focus on the multi-agent system, proposing the Permutation-Invariant Adversarial Attack Method. It models the attack path as the Maximum-Flow Minimum-Cost Problem, and combines the Permutation-Invariant Evasion Loss to optimize prompt propagation, improving the attack success rate by up to seven times. These examples underscore the critical threat of cross-agent contamination. To better understand the vulnerabilities of multi-agent systems, we examine the key attack types of attacks in detail.
R2: Privacy Leakage. The communication process with multiple agents will suffer from the risk of information leakage. Different from the user-agent interaction, such leakage is conducted by agents instead of users. Besides, these kinds of attacks include both the malicious sniffing or stealing of sensitive information and the inadvertent information spreading from high-authority agents to low-authority agents. We think the latter may be more difficult to detect. Kim et al. [kim2025prompt] show that, in permission escalation attacks, malicious agents can generate adversarial prompts or inject unsafe data to cause unauthorized attacks.
R3: Description Poisoning. Multi-agent systems rely heavily on agents’ capability descriptions, role definitions, and self-reported functional metadata to determine how agents should be routed, composed, or delegated. Description poisoning occurs when an agent’s semantic descriptors are tempered by embedding misleading role definitions, fabricated capabilities, or covert prompt instructions. Such poisoning manipulates how the system interprets the agent’s purpose, leading to incorrect routing decisions, assigning inappropriate permissions, or over-trusting a malicious agent [soloio2024mcp_a2a_attacks, narajala2025securing].
(ii) Unique Risks in CS-based Communication. In client–server architectures, the server acts as the sole semantic arbiter responsible for interpreting agent capabilities, normalizing intermediate semantics, and orchestrating cross-agent routing. This centralized semantic authority introduces several unique L3 risks.
R4: Centralized Poisoning. In CS-based multi-agent systems, the server is responsible for interpreting agent roles, capabilities, and routing rules. If an attacker compromises this centralized semantic state through prompt injection, metadata forgery, or poisoning of the server-side LLM, a malicious agent may be falsely labeled as a trusted safety auditor or privileged tool agent. Because all agents rely on server-maintained semantics, the misclassification becomes systemic, influencing every downstream interaction.
R5: Semantic Rewriting Attack. Most CS-based orchestrators normalize or rewrite user queries and intermediate agent outputs before forwarding them. If the server is compromised or jailbroken, it can subtly modify task intent, safety constraints, or execution semantics during this rewriting step while preserving superficially valid syntax. This creates a semantic attack channel unique to centralized architectures that enables covert manipulation of task objectives.
(iii) Unique Risks in P2P-based Communication. In a P2P architecture, the lack of a centralized validation and arbitration mechanism makes some risks particularly severe.
R6: Cognitive and Ethical Drift. In decentralized P2P architectures, agent heterogeneity introduces a fundamental vulnerability: divergence in semantic interpretation. This risk manifests in a two-stage cascade. At the cognitive layer, the system’s semantic interoperability is compromised. Agents powered by disparate LLM backbones (e.g., different models, versions, or fine-tuning protocols) will inevitably exhibit varied interpretations in the same context, which could trigger conflicting actions. This semantic ambiguity guarantees desynchronized collaboration and mission degradation. More critically, this cognitive divergence predictably escalates to the normative layer. When agents embodying different ethical frameworks or value systems confront a dilemma, such as scarce resource allocation, they arrive at mutually exclusive moral judgments. Without a central arbiter to enforce a global ethical policy, a persuasive but misaligned agent can hijack the collective decision-making process, causing a systemic Ethical Drift. This escalation from cognitive misunderstanding to normative collapse is not merely a risk, but a fundamental challenge to the viability of decentralized autonomous systems.
R7: Contextual Fragmentation. P2P communication architectures, characterized by multi-hop, asynchronous message passing, inherently lead to contextual fragmentation. Unlike CS systems, each P2P agent maintains only a localized and potentially inconsistent perception of the global conversational state. This compels agents to operate on incomplete or stale information, forcing them to perform erroneous state inference to bridge contextual gaps. Consequently, actions may be based on outdated premises, and the nuances of complex, multi-turn dialogues can be lost, leading to profound misinterpretations of intent. This is not merely a data consistency issue; it is a breakdown of shared understanding at the semantic level, resulting in action desynchronization and the failure to achieve coherent joint goals.